# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathrm{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
@@
Now we start from (A.3). In the case of different means and same variance, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-e_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{1}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$-\frac{1}{2N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
@@
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}+(\hat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)$
@@
by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Thus, (A.15) is verified. The form of (A.15) can be further refined under $H_{0}$, $H_{1}$ and $H_{1L}$ respectively. For example, under $H_{0}$, we can write $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4. $\square$
Proof of Theorem 7.
The technique is still the same as the previous ones. Recall that in the context of same mean and different variances, we denote $p_{j}=N_{j}/N$, $q_{j}=JN_{j}/N$ for $1\leq j\leq J$, and we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
We first tackle the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$. For $\hat{\mu}-\mu$, note that $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here. Then,
$\hat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
From (A.20) we obtain $\mathrm{E}[\hat{\mu}-\mu]=0$, and $\operatorname{Var}[\hat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$. Then $\hat{\mu}-\mu=O_{p}(N^{-1/2})$. For $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
$\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
@@
by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\to_{d}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\to_{d}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
@@
$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
@@
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
@@
$\mathrm{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
@@
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
@@
Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}$
$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}.$
@@
$D_{N1}$ $=-\frac{\hat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$

### v4

@@ A.4 Proof of Theorem 4, 5 and 6
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathbb{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathbb{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathbb{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
@@
Now we start from (A.3). For model (3.9), (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-e_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}$
$=$ $-\frac{1}{\widehat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$-\frac{\widehat{\sigma}^{2}-\sigma^{2}}{\widehat{\sigma}(\widehat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\widehat{\sigma}^{2}-\sigma^{2}\right)$
$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\widehat{\sigma}^{2}-\sigma^{2}\right)$
$=$ $-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$-\frac{1}{2N}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
@@
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{2}$
$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\widehat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\widehat{\sigma}^{2}-\sigma^{2})}{\widehat{\sigma}\sigma(\widehat{\sigma}+\sigma)}\right)^{4}\right]$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}+(\widehat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathbb{E}\left[e^{4}\right]+o_{p}(1)\right)$
@@
by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\widetilde{\varepsilon}_{ij}^{2}/\widetilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\widetilde{\varepsilon}_{ij}^{2}}{\widetilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Thus, (A.15) is verified. The form of (A.15) can be further refined under $H_{0}$, $H_{1}$ and $H_{1L}$ respectively. For example, under $H_{0}$, we can write $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4. $\square$
A.5 Proof of Theorem 7
The technical arguments remain the same as in the previous cases. Recall that $p_{j}=N_{j}/N$ and $q_{j}=JN_{j}/N$ for $1\leq j\leq J$. And for model (3.10), we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$ (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
We first tackle the estimation effects of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}$. For $\widehat{\mu}-\mu$, note that $\widehat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here. Then,
$\widehat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
From (A.20) we obtain $\mathbb{E}[\widehat{\mu}-\mu]=0$, and $\operatorname{Var}[\widehat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$. Then $\widehat{\mu}-\mu=O_{p}(N^{-1/2})$. For $\widehat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
$\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\widehat{\mu}-\mu\right)^{2}-2\left(\widehat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
@@ A.5 Proof of Theorem 7
by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\overset{d}{\to}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\overset{d}{\to}\mathcal{N}(0,\mathbb{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
@@
$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
@@
$=$ $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
@@
$\mathbb{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
@@
$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathbb{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
@@
Now we start from (A.3). For model (3.10), (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\widehat{\varepsilon}_{ij}}{\widehat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\widehat{\varepsilon}_{ij}-\varepsilon_{ij}}{\widehat{\sigma}_{j}}$
$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\widehat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\widehat{\sigma}_{j}\sigma_{j}(\widehat{\sigma}_{j}+\sigma_{j})}.$
@@
$D_{N1}$ $=-\frac{\widehat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$